Exercise 5-18 Break-Even and Target Profit Analysis; Margin of Safety; CM Ratio (LO5-1, LO5-3, LO5-5, LO5-6, LO5-7) Menlo ComReq 1 Req 2 Req ЗА Req 3B Req 4 Req 5 What is the monthly break-even point in unit sales and in dollar sales? Break-even poinReq 1 Req 2 Req ЗА Req 3B Req 4 Req 5 Without resorting to computations, what is the total contribution margin at the break-eReg1 Req 1 Req 2 Rega | Req3A | Req3B Req ЗА Req 3B Rega Req5 Req 4 Req 5 How many units would have to be sold each month toReq 1 Req 2 Req ЗА Req 3B Req 4 Req 5 Verify your answer by preparing a contribution format income statement at the target saRega Rega Rea Sa Reg 38 Req 1 Rega Reqs Req 2 Req ЗА Req 3B Req 4 Req 5 Refer to the original data. Compute the companys marReq 1 Req 2 Req ЗА Req 3B Req 4 Req 5 What is the companys CM ratio? If sales increase by $50,000 per month and there is no
Answer #1

1) Break even unit = Fixed cost/Contribution margin per unit = 216000/18 = 12000 Units

Break even sales = 12000*30 = $360000

2) Contribution margin = 216000

3a) Required unit sales = (216000+90000)/18 = 17000 Units

3b) Contribution margin income statement

Total Per unit
Sales 510000 30
Variable cost 204000 12
Contribution margin 306000 18
Fixed cost 216000
Net income 90000

4) Calculate following

Dollar Percentage
Margin of safety 450000-360000 = 90000 90000/450000 = 20%

5) CM ratio = 18/30 = 60%

Net operating income increase by = 50000*60% = 30000